Bondi Corporation makes automotive engines. For the most recent month, budgeted production was 1,800 engines. The standard power cost is $3.40 per machine-hour. The company's standards indicate that each engine requires 9.6 machine-hours. Actual production was 2,100 engines. Actual machine-hours were 19,290 machine-hours. Actual power cost totaled $69,110.

Required:

Determine the rate and efficiency variances for the variable overhead item power cost. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

Rate Variance = (SR-AR) * AMH

= (3.40 - 69,110/19,290)*19,290

= 3524 U

Efficiency variance = (SH-AH) * SR

= (9.6*2100 - 19,290)*3.40

= 2958 F